One electron and one proton of a hydrogen atom are at a distance of 2.2✕10−11m apart. What is the potential energy, in joules, of the electron in the proton’s field?






1
Expert's answer
2022-03-04T08:30:48-0500

The electrostatic potential energy

"E_p=k\frac{q_1q_2}{r}"

"E_p=9*10^9*\frac{-1.6*10^{-19}*1.6*10^{-19}}{2.2*10^{-11}}=-1.05*10^{-17}\:\rm J"